Two waves travel in opposite directions Due in 7 hours, 55 minutes A generator at one end of a very long string creates a wave given by Yix,t) -(2.00 cm)cost(n/2)(6.47x+ 9.34s1)and a generator at the other end creates the wave y2(x,t) (2.00 cm)*cos[(n/2)*(6.47*x - 9.34 s 1 *t)], where x is in meters and t is in seconds for both waves. Calculate the frequency of each wave Submit Answer Incorrect. Tries 4/99 Previous Tries Calculate the period of each wave. Submit Answer Incorrect. Tries 1/99 Previous Tries Calculate the wavelength of each wave. Submit Answer Tries 0/99 Calculate the speed of each wave. Submit Answer Tries o/99 For x20, what is the location of the node having the smallest value of x? Assume the first antinode is where the string is being driven, x 0 Submit Answer Tries 0/99 For x20, what is the location of the node having the second smallest value of x? Submit Answer Tries 0/99 For x20, what is the location of the node having the third smallest value of x? Submit Answer Tries o/99 For x20, what is the location of the antinode having the smallest value of x? Submit Answer Tries /99 For x20, what is the location of the antinode having the second smallest value of x? Submit Answer Tries 0/99 For x20, what is the location of the antinode having the third smallest value of x? Submit Answer Tries 0/99
Answer #1

Answer :

If you write y = A cos 2π(x/L+ft) or y = A cos(kx +wt)

where x is distance

t is time

L is wavelength

and f is frequency

Wave speed is f*L.

Compare with the given equation :

y = (2.00 cm) cos[(π/2)(6.47 m-1x - 9.34 s-1t)]

multiply π/2 to inside equation and get

y = (2.00 cm) cos[(π x 3.235 m-1x - 4.67 x π s-1t)]

1) frequency = w/2π

= 4.67 x π/ 2π= 2.335 s-1

2) wavelength = 2π/k = 2π/3.235π = 0.618 m

3) speed of the wave = f x w = 2.335 x 0.618 = 1.44 m/s